OpenStudy (anonymous):

3lnx-ln(x+1)=ln(x-1) Solve for x

OpenStudy (bahrom7893):

ln(x^3/(x+1))=ln(x-1)

OpenStudy (bahrom7893):

x^3/(x+1)=(x-1)

OpenStudy (bahrom7893):

x^3=x^2-1

OpenStudy (bahrom7893):

x^3-x^2+1=0

OpenStudy (bahrom7893):

hmm im lost now..

OpenStudy (anonymous):

Solve with quadratic formula?

OpenStudy (anonymous):

not a quadratic...

OpenStudy (anonymous):

Oh wow i'm an idiot.

OpenStudy (mertsj):

It's not quadratic. If there are any rational roots they are factors of 1. So use synthetic division with 1 and/or -1 on the "shelf" and see if either is a root.

OpenStudy (mertsj):